OpenStudy (anonymous):

x^2 + y^2 - 8X + 12Y - 20 = 0 What the radius?

OpenStudy (anonymous):

x^2-8x+y^2+12y=20 x^2-8x+(16)+y^2+12y+36=20+16+36 (x-4)^2+(y+6)^2=72 r^2=72 \[r= \sqrt{72}\approx8.4\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

by complete the square for x & y.

OpenStudy (anonymous):

did you use the equation (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

yes